OpenStudy (anonymous):

@AlexandervonHumboldt2 hiiiii!! :) I need help!

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

\[\frac{ 5 }{ -8}+\frac{ 5 }{ 3}\]

OpenStudy (alexandervonhumboldt2):

ok so 5/(-8)=-5/8

OpenStudy (alexandervonhumboldt2):

multiply first fractor by 3/3. tell me what you get

OpenStudy (alexandervonhumboldt2):

fraction i meant

OpenStudy (alexandervonhumboldt2):

you neeed to make demonominators the same

OpenStudy (anonymous):

the denominator will be 24

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (alexandervonhumboldt2):

so you have -15/24 + 40/24

OpenStudy (alexandervonhumboldt2):

=??

OpenStudy (anonymous):

i got 25/24

OpenStudy (alexandervonhumboldt2):

yes 25/24 is correct

OpenStudy (anonymous):

yay! awesome so final answer don't have to simplify right?

OpenStudy (anonymous):

@sammixboo hiiiiii :)

sammixboo (sammixboo):

Hola

sammixboo (sammixboo):

Yes, that problem can be turned into a mixed fraction. DO you know how to do that?

OpenStudy (anonymous):

hi bestie haha! is my answer is right i need help turning it to a mixed fraction :(

OpenStudy (anonymous):

@sammixboo

sammixboo (sammixboo):

Ok, so how many times does 24 go into 25?

OpenStudy (anonymous):

uh let me does it go in one time right?

sammixboo (sammixboo):

Right! So our big number ( whole number is 1 ) and we know our denominator is going to be 24, so that means we have this! \(1\dfrac{?}{24}\)

sammixboo (sammixboo):

To find what does in the question mark, just subtract 25 - 24

OpenStudy (anonymous):

so its going to be \[1\frac{ 1 }{ 24 }\]

OpenStudy (alexandervonhumboldt2):

yes

OpenStudy (anonymous):

@AlexandervonHumboldt2 thank you :)